OpenStudy (anonymous):

int((2xdx)/sqrt(2x^2 - 2x +5)) I have problem in finding a solution for this one. I understand that i have to use substitution.

OpenStudy (perl):

youre going to want to complete the square in the denominator

OpenStudy (anonymous):

\[\int\limits\frac{ 2xdx }{ \sqrt{ 2x^2 -4x +5 } } = \int\limits\frac{dx^2}{\sqrt{2x^2 -4x +5}}=\] As i understood.

OpenStudy (loser66):

x cannot be "time" to dx to get dx^2 nope. as perl suggested, just complete the square in the denominator first

OpenStudy (loser66):

@perl back here to finish your starting.

OpenStudy (perl):

ok we can complete the square

OpenStudy (perl):

this other problem is driving me nuts

OpenStudy (anonymous):

\[\int\limits\frac{2xdx}{\sqrt{2(x^2 -2x +2\frac{ 1 }{ 2 } +1\frac{1}{2}-1\frac{1}{2})}}=\int\limits\frac{2xdx}{\sqrt{2(x^2 -2^2-1\frac{1}{2})}}=\]

OpenStudy (perl):

ok

OpenStudy (perl):